OpenStudy (anonymous):

Find the first three terms of the arithmetic sequence in which n=39, a of n = 134.4, and S of n = 5538

OpenStudy (campbell_st):

ok... so you are given n = 39 so for the 39th term you have \[134.4 = a + 39 - 1) \times d.....or...... 134.4 = a + 38d\] now for the sum of 39 terms \[5538 = \frac{39}{2} \times [ 2a + (39 - 1) d] ... or.... 5538 = \frac{39}{2}[2a + 38d]\] simplify the 2nd equation to 11076 = 78a + 1482d So I would use substitution take the 1st equation and make a the subject a = 134.4 - 38d and substitute it into the 2nd equation 11076 = 38(134.4 - 38d) + 1482d 11076 = 5107.2 - 1444d + 1482d solve for d. when you get d, substitute it into either of the equations to find a, the 1st term. hope it helps

OpenStudy (anonymous):

Thank you so much, this really helps